# Adversarial Agent Collaboration for C to Rust Translation

### v1

@@ Abstract
Translating C to memory-safe languages, like Rust, prevents critical memory safety vulnerabilities that are prevalent in legacy C software. Existing approaches for C to safe Rust translation, including LLM-assisted ones, do not generalize on larger (>500>500 LoC) C codebases because they depend on complex program analyses that frequently break. In this work, we present ACToR (Adversarial C To Rust translator), a simple LLM agent-based approach. Inspired by GANs, ACToR pits a generator agent against a discriminator agent, which collaborate to iteratively generate a Rust translation. On each iteration, the translator agent synthesizes and refines a Rust translation to pass an existing suite of tests, and then the discriminator agent finds new failing tests. We demonstrate that ACToR translates all of the 63 real-world command line utilities considered in our benchmarks, which have an average size of 485 lines of code, and it achieves over 90% test pass rate with zero human intervention. To our knowledge, it is the first such system that reliably translates C programs of this scale. Furthermore, ACToR improves translation correctness by up to 18.9% compared to baseline, non-adversarial approaches.
@@ 1 Introduction
Memory safety vulnerabilities in C/C++ code constitute a large fraction of security vulnerabilities each year. Microsoft reported 70% of their CVEs are due to memory safety issues (Microsoft, 2019a; b). In response, there has been an increasing demand to translate memory-unsafe legacy code to modern memory-safe languages like Rust from industry and governments (DARPA, 2025), thereby ensuring full memory safety by design. Unfortunately, given the scale of legacy code that permeates our software systems today, manual translation of (millions of lines of) existing C code is a formidable and practically infeasible task. A reasonable translation from C to Rust must use safe and idiomatic Rust programming abstractions, which are checked mostly at compile time and avoid runtime performance overhead, unlike memory safety checking directly in C (Nagarakatte et al., 2010; 2009; Orthen et al., 2024; Serebryany et al., 2012).
Existing approaches for automated translation come in two flavors: rule-based and LLM-assisted. Rule-based translators, like c2rust (Immunant, ) and its derivatives (Zhang et al., 2023; Emre et al., 2023; Ling et al., 2022), produce unidiomatic and unsafe Rust, which limits their usefulness in practice. LLM-assisted translators produce more idiomatic and safe Rust; however, current works along this line fail to generalize to large (>500>500 lines of code) C projects. No prior work on LLM-assisted translation has demonstrated their approach successfully translating more than one large C project, and several reports require manual intervention owing to this brittleness (Shetty et al., 2024; Cai et al., 2025). This brittleness is largely due to their dependency on complex program analyses that break on unseen C programs.
@@
While the ability to access development environment without a priori fixed workflow is a promising first step towards more reliable and realistic translations, we find, rather unsurprisingly, that simplistic code agents like the one described above nevertheless produce incorrect translations; the behavior of the translated Rust code they produce diverges significantly from that of the original C code on inputs absent from the test suite. As a result, such a vanilla agentic setup requires frequent developer intervention and offers little to no improvement over traditional LLM-based translation techniques. At a high level, a simplistic agentic setup, like the one we just described, lacks the key ability of generalization — it produces a translation that is overfitted to the (typically small) test suite available to it, but often cannot ensure expected behavior on inputs absent from the suite. Unfortunately, as our evaluation shows, simple fixes such as by increasing the test coverage on the C source, which we achieve by augmenting agents with standard test coverage tools (GNU, 2025), continue to remain ineffective in enhancing the quality of translations.
@@
We demonstrate the effectiveness of ACToR on two sets of benchmarks: A micro-benchmark set of $6$ C programs considered in recent work, and, a much larger macro-benchmark set of $57$ C programs from the standard BSDCoreUtils. These $63$ real-world C programs range from $7$ to $5,469$ lines of code (LoC), with a median of 485, and totaling over 30,000. ACToR automatically produces Rust translations for all benchmarks that pass more than 90% of tests on average. To our knowledge, it is the first system to yield translations with zero human intervention and with high pass rates for this scale of programs. In addition, we show that ACToR’s adversarial design choice improves pass rates by up to 18.9% compared with the baseline, and it also improves the quality of tests generated, as measured by their ability to find discriminating behavior.
@@ 2 Related Work
Automated translation of C to Rust is a relatively recent approach towards achieving full memory safety (Emre et al., 2021). The desired goal of such translation is to generate safe [^1] The use of unsafe keyword in Rust allows mixing unchecked and possibly memory unsafe code in otherwise safe Rust Rust code corresponding C code, i.e., Rust code that uses only idiomatic Rust primitives that the compiler can statically verify or trust to be memory safe. While several prior works have tried, none have worked reliably even for programs of moderate size (e.g., about $500$ LoC) for varied reasons (Li et al., 2024).
@@ 2.1 Program Translation (C to Rust)
Existing works on C-to-Rust translation can primarily be categorized as rule-based or LLM-based techniques. Rule-based techniques such as c2rust (Immunant, ) and corrode (Sharp, ) opt for mechanical, line-by-line, rewriting of the C code, often resorting to raw pointers and unsafe Rust. Later efforts explored ways to infer Rust types from C code, either by compiler feedback (Emre et al., 2023; Hong & Ryu, 2025b), SMT solving (Zhang et al., 2023), data-flow graphs (Xu & Huang, 2025), or syntactic rewriting (Ling et al., 2022), and specialized analyses targeted idioms like tagged unions (Hong & Ryu, 2024b), output parameters (Hong & Ryu, 2024a), and file APIs (Hong & Ryu, 2025a). Beyond general C, Low* compilation to Rust (Fromherz & Protzenko, 2024) demonstrates scalable translation in restricted settings, though such techniques remain infeasible for general-purpose C due to expressiveness and unmodeled APIs.
@@ 2.2 Collaborative LLMs & Agents
Many works propose multi-LLM and multi-agent systems, where the LLMs or agents collaborate to produce a final answer. Many works on code generation propose specific roles, such as coder, tester, or analyst, which collaborate to write code (McAleese et al., 2024; Chen et al., 2023; Dong et al., 2024; Huang et al., 2023; Qian et al., 2024; Islam et al., 2024). Collaborative approaches have been used outside of code generation as well. Many works propose self-reflection techniques (Shinn et al., 2023; Madaan et al., 2023; Dhuliawala et al., 2023), sometimes aided by tools (pan, 2024; Gou et al., 2024; Dong et al., 2025), where the LLM first proposes an answer, then reflects on its answer to improve accuracy. Similar works propose multi-agent debate architectures (Chan et al., 2024; Du et al., 2023; Liang et al., 2023; Hong et al., 2024; Wu et al., 2024; Liu et al., 2024) where multiple answers are proposed, then the LLM(s) debate the correct answer. Several works propose frameworks for agent collaboration (Hong et al., 2024; Wu et al., 2024; Liu et al., 2024).
The key difference between ours and prior work is in the high-level principle: To reliably improve the translated code, we leverage the C program as the oracle in the construction of a discriminator, which automatically identifies correctness gaps and provides iterative feedback to the translator.
@@ 3.1 Code Translation Task
Given a source C program $c$, the goal of the translation is to find a functionally correct Rust code $rs$ that has equivalent external behaviors as the source C program on all valid inputs. External behaviors include standard output, error messages, file modifications, and so on. More specifically, we want the translated code to satisfy the boolean requirement $\texttt{IsEq}(c,rs)$, where IsEq is an oracle that takes a C program and a Rust program as input, and determines whether they have equivalent behaviors on the valid universe $U$ of inputs that do not violate memory safety. This objective naturally captures the desired requirement that the translated Rust program has the same behavior as the C program on safe inputs. On the other hand, memory-safe handling of malicious inputs is guaranteed by the Rust compiler itself.
A naive approach to obtaining a Rust translation is to prompt the LLM agent with the C code $c$ and a small finite set $T\subset U$ of inputs sampled from the safe input universe $U$ with the goal of computing a Rust program that matches the behaviour of $c$ on all inputs from the sampled set $T$. Unfortunately, such a naive method does not produce a satisfactory translation. The space of valid inputs is typically infinite for most programs. Indeed, we empirically observed that with only a limited number of input instances tested one-shot by the agents, the produced translation does not generalize to unseen test cases beyond the limited set $T$.
We remark that a perfect oracle IsEq does not exist in practice. There is no automated tool that practically works in determining whether the given C code and Rust are equivalent in the input universe, or provides counterexamples when they mismatch. On the contrary, what we have is ‘IsEq*’, which takes as input both the C and the Rust code, together with a concrete input $t$, tells whether they mismatch on this input $t$. This intuitively means we run the C code and Rust code on the same input, and check if they result in the same effect. The problem now becomes how to satisfy IsEq with only IsEq* available.
@@ 3.2 Adversarial Translation
4:while Not PassesSuite($c,rs,T_{0}$) do $\triangleright$ initial translation on seed tests
@@
8:for $k=1$ to MaxIter do
9:  $numRetry\leftarrow 0$
10:  repeat
11:   $Batch\leftarrow$ Discriminator($c,rs,n=\text{TestBatchSize}$)
12:   $IsValid\leftarrow\big(\forall t\in Batch:\;\textsc{ValidTestFormat}(t)\land\textsc{SanityOk}(t,c)\big)$
13:   $IsDisc\leftarrow\big(\exists t\in Batch:\;\neg\texttt{IsEq}^{*}(c,rs,t)\big)$
14:   $numRetry\leftarrow numRetry+1$
15:   $IsReachingLimit\leftarrow(numRetry\geq\text{RetryLimit})$
16:  until $(|Batch|=\text{TestBatchSize}~\land~IsValid~\land~IsDisc)~\lor~IsReachingLimit$
17:  $T\leftarrow T\cup Batch$
18:  $numRetry\leftarrow 0$
19:  while (Not PassesSuite($c,rs,T$)) $\lor~numRetry\leq\text{RetryLimit}$ do
20:   $numRetry\leftarrow numRetry+1$
21:   $rs\leftarrow$ Translator($c,rs,T$)
22:  end while
23:end for
24:return rsrs as $rs^{*}$, $T$ as $T^{*}$
ACToR takes as input the C code $c$, the universe $U$, the seed test set $T_{0}$, together with some iteration parameters such as the number of times the translator and discriminator interact (MaxIter) before ACToR outputs the final translated program. ACToR then conducts iterative translation on the source C code and outputs both the Rust translation rsrs* and the final test set TT*. During iterative improvement, ACToR maintains an append-only test set $T$. In the beginning, the translator agent generates the initial translation and ensures it passes all the seed tests (lines 1-3) before the adversarial iterative loop (involving the discriminator) begins. Next, in each iteration that follows, the discriminator agent reads the C code $c$ and the current Rust code $rs$, and generates test cases on which C and Rust outputs disagree (lines 7-13). The discriminator is structured so that (1) all the test cases abide by the correct format and also pass basic sanity checks when running them on the source C program (line 9); and that (2) some test cases witness a mismatch between outputs of the C and the Rust code that the translator has created so far (line 10). We ensure this by providing a prompt to the discriminator to continue retrying until the generated test cases meet the above two requirements. Once successful, the newly generated test cases ($Batch$) are added to the full test set. The translator agent then takes the updated test cases as guidance to improve the translation. It is set up with prompts that repeatedly ask it to pass all the test cases in the current test set (lines 16-19). If the translator runs out of retries but still cannot fully pass the test set $T$, we let it skip and move to the next iteration. After a fixed maximum number of iterations, ACToR stops the loop and outputs the final translation and the test set generated.
@@ 4 Evaluation
The goal of a good defense is to retain the functionality of the given C code, and make the defended program fully memory-safe. Specifically, we evaluate ACToR on the following criteria:
@@
    Ablation. How much does the adversarial agentic design contribute to correctness?
@@ 4.1 Experiment Settings
Our main macro-benchmark is a set of 57222The full set of BSDCoreUtils contains 69 programs, but we filter out 12 of them from our main results due to three reasons: (1) some utilities are trivial (e.g., true, yes) (2) certain utilities require special environments (e.g., stty) or privileges (e.g., sudo) not available in our experimental sandbox; and (3) certain utilities perform destructive operations (e.g., rmdir) that may break the testing harness. real-world system utility programs (e.g., ls, cp, xargs) obtained from a software project named BSDCoreUtils. It has an average of 492 lines of code each. We use this benchmark to evaluate the scalability of ACToR, and compare correctness to our baselines. These programs do not come with tests or test runners. We preprocessed these programs to add test runners, which can automatically run a set of given tests for each program, compare C and Rust outputs, as well as produce coverage reports. Note that while we provide the test runner, it is infeasible to manually create comprehensive test cases for all programs to evaluate which translation strategy is more correct. Thus, unlike micro-benchmarks, we perform a relative evaluation between ACToR and baselines, as described below in the Evaluation Metrics section.
To minimize measurement noise and make our experiments automatable, our evaluation focuses on single-threaded C programs that have deterministic behaviors across different executions, which is currently an assumption of our testing framework that both the translator and the discriminator depend on. Table 1(b) summarizes the statistics of two benchmarks, where we categorize the 57 programs into five categories: process execution control, file/directory management, shell scripting commands, system information access, and textural content processing.
Agent Frameworks and Models. We use Claude Code (Anthropic, 2025a) with Claude Sonnet 4 (Anthropic, 2025b) as the main agent framework and LLM for our experiments. To test if our approach generalizes to other agent frameworks and LLMs, we also performed experiments using Mini-SWE-Agent (SWE-agent, 2025), a popular open-source agent framework, with both GPT 5 mini (OpenAI, 2025) and Claude 4 Sonnet as the LLM. We use ‘ClaudeCode’, ‘SWE+Sonnet’, and ‘SWE+GPT5mini’ to represent them in later sections, respectively.
@@
Evaluation Metrics. Our primary evaluation metric to evaluate the correctness of an individual translation is pass rate, which is the percentage of tests that pass for a given set of tests. For our micro-benchmark, we use our manually written test cases to compute pass rates for each benchmark. For our macro-benchmarks, which do not have manually written tests, we compute pass rates using a set of tests produced by a competing method, and refer to this metric as relative pass rate. For example, when comparing ACToR with the coverage baseline, we compute pass rate for ACToR using the tests generated by the coverage baseline, and vice versa. We use this approach because a given method might get biased towards its own tests (usually a given method passes all of its own tests). Relative pass rate provides a strong signal for comparing ACToR against our baselines. In addition, the test sets used to compute relative pass rate achieve an average of 89.7% line coverage on the micro benchmark and 90.6% line coverage on the macro benchmark, by unioning test cases across translation methods. This indicates they are a useful signal for overall correctness as well.
Experimental Configuration. ACToR and all baselines start with 15 manually crafted, diverse seed tests based on the C code for each program  [^3] The seed tests carry intended test format that both translator and discriminator of ACToR use. This avoids a cold-start for ACToR as well as all other systems we compare to.. For the coverage baseline, ACToR-NoFuzz, and ACToR, we run 10 iterations of test generation and translation. On each iteration, we generate 3 new tests, resulting in a total of 45 tests per program in the end. We choose 10 iterations because more iterations usually do not result in improvements in pass rate and limit the incurred financial costs. We generate 3 tests per iteration because larger values lead to a decline in quality of tests, and smaller values have slower convergence. In considering cost, we allow translator agents and discriminator agents to proceed directly to the next iteration if they are unable to make progress (i.e., generate valid test cases or correct the translation to pass the current tests) after 3 retries. The prompt for the translator agent is shared among all the translation methods and put in Appendix B. The prompts for the test generator agent in the coverage baseline and the discriminator of ACToR are in Appendix B.

### v2

@@ Abstract
Translating C to memory-safe languages, like Rust, prevents critical memory safety vulnerabilities that are prevalent in legacy C software. Existing approaches for C to safe Rust translation, including LLM-assisted ones, do not generalize on larger (>500>500 LoC) C codebases because they depend on complex program analyses that frequently break. In this work, we present ACToR (Adversarial C To Rust translator), a simple LLM agent-based approach. Inspired by GANs, ACToR pits a generator agent against a discriminator agent, which collaborate to iteratively generate a Rust translation. On each iteration, the translator agent synthesizes and refines a Rust translation to pass an existing suite of tests, and then the discriminator agent finds new failing tests. We demonstrate that ACToR translates all of the 63 real-world command-line utilities considered in our benchmarks, which have an average size of 473 lines of code, and it achieves over 90% test pass rate with zero human intervention during translation. To our knowledge, it is the first work to show evidence that an agent-centric approach can reliably and automatically convert standalone command-line C programs at this scale. Furthermore, ACToR improves translation correctness by up to 25.1% compared to baseline, non-adversarial approaches.
@@ 1 Introduction
Memory safety vulnerabilities in C/C++ code constitue a large fraction of security vulnerabilities each year. Microsoft reported 70% of their CVEs are due to memory safety issues (Microsoft, 2019a; b). In response, there has been an increasing demand to translate memory-unsafe legacy code to modern memory-safe languages like Rust from industry and governments (DARPA, 2025), thereby ensuring full memory safety by design. Unfortunately, given the scale of legacy code that permeates our software systems today, manual translation of (millions of lines of) existing C code is a formidable and practically infeasible task. A reasonable translation from C to Rust must use safe and idiomatic Rust programming abstractions, which are checked mostly at compile time and avoid runtime performance overhead, unlike memory safety checking directly in C (Nagarakatte et al., 2010; 2009; Orthen et al., 2024; Serebryany et al., 2012).
Existing approaches for automated translation come in two flavors: rule-based and LLM-assisted. Rule-based translators, like c2rust (Immunant, ) and its derivatives (Zhang et al., 2023; Emre et al., 2023; Ling et al., 2022), produce unidiomatic and unsafe rust, which limits their usefulness in practice. LLM-assisted translators produce more idiomatic and safe rust, however current works along this line fail to generalize to large (>500>500 lines of code) C projects. No prior work on LLM-assisted translation has demonstrated their approach successfully translating more than one large C project, and several report requiring manual intervention owing to this brittleness (Shetty et al., 2024; Cai et al., 2025). This brittleness is largely due to their dependency on complex program analyses that break on unseen C programs.
@@
While the ability to access development environment, without an a priori fixed workflow is a promising first step towards more reliable and realistic translations, we find, rather unsurprisingly, that simplistic code agents like the one described above nevertheless produce incorrect translations; the behavior of the translated Rust code they produce diverges significantly from that of the original C code on inputs absent from the test suite. As a result, such a vanilla agentic setup requires frequent developer intervention and offers little to no improvement over traditional LLM-based translation techniques. At a high level, a simplistic agentic setup, like the one we just described, lacks the key ability of generalization — it produces a translation that is overfitted to the (typically small) test suite available to it, but often cannot ensure expected behavior on inputs absent from the suite. Unfortunately, as our evaluation shows, simple fixes such as by increasing the test coverage on the C source, which we achieve by augmenting agents with standard test coverage tools (GNU, 2025), continue to remain ineffective in enhancing the quality of translations.
@@
We demonstrate the effectiveness of ACToR on two sets of benchmarks: A micro-benchmark set of $6$ C programs considered in recent work, and, a much larger macro-benchmark set of $57$ C programs from the standard BSDCoreUtils. These $63$ real-world C programs range from $7$ to $5,469$ lines of code (LoC), with a median of 473, and totaling around 30,000. ACToR automatically produces Rust translations for all benchmarks that pass more than 90% of tests on average with zero human intervention during translation. To our knowledge, it is the first work to show evidence that an agent-centric approach can reliably and automatically convert standalone command-line C programs at this scale. In addition, we show that ACToR’s adversarial design choice improves pass rates by up to 25.1% compared with the baseline, and it also improves the quality of tests generated, as measured by their ability to find discriminating behavior.
@@ 2 Related Work
Automated translation of C to Rust is a relatively recent approach towards achieving full memory safety (Emre et al., 2021). The desired goal of such translation is to generate safe [^1] The use of unsafe keyword in Rust allows mixing unchecked and possibly memory unsafe code in otherwise safe Rust Rust code corresponding C code, i.e., Rust code that uses only idiomatic Rust primitives that the compiler can statically verify or trust to be memory safe. While several prior works have tried, none have worked reliably even for programs of moderate size (e.g. about $500$ LoC) for varied reasons (Li et al., 2024).
@@ 2.1 Program Translation (C to Rust)
Existing works on C-to-Rust translation can primarily be categorized as rule-based or LLM-based techniques. Rule-based techniques such as c2rust (Immunant, ) and corrode (Sharp, ) opt for mechanical, line by line, rewriting of the C code, often resorting to raw pointers and unsafe Rust. Later efforts explored ways to infer Rust types from C code, either by compiler feedback (Emre et al., 2023; Hong & Ryu, 2025b), SMT solving (Zhang et al., 2023), data-flow graphs (Xu & Huang, 2025), or syntactic rewriting (Ling et al., 2022), and specialized analyses targeted idioms like tagged unions (Hong & Ryu, 2024b), output parameters (Hong & Ryu, 2024a), and file APIs (Hong & Ryu, 2025a). Beyond general C, Low* compilation to Rust (Fromherz & Protzenko, 2024) demonstrates scalable translation in restricted settings, though such techniques remain infeasible for general-purpose C due to expressiveness and unmodeled APIs.
@@ 2.2 Collaborative LLMs & Agents
Many works propose mutli-LLM and multi-agent systems, where the LLMs or agents collaborate to produce a final answer. Many works on code generation propose specific roles, such as coder, tester, or analyst, which collaborate to write code (McAleese et al., 2024; Chen et al., 2023; Dong et al., 2024; Huang et al., 2023; Qian et al., 2024; Islam et al., 2024). Collaborative approaches have been used outside of code generation as well. Many works propose self-reflection techniques (Shinn et al., 2023; Madaan et al., 2023; Dhuliawala et al., 2023), sometimes adied by tools (pan, 2024; Gou et al., 2024; Dong et al., 2025), where the LLM first proposes an answer, then reflects on its answer to improve accuracy. Similar works propose multi-agent debate architectures (Chan et al., 2024; Du et al., 2023; Liang et al., 2023; Hong et al., 2024; Wu et al., 2024; Liu et al., 2024) where mutliple answers are proposed, then the LLM(s) debate the correct answer. Several works propose frameworks for agent collaboration (Hong et al., 2024; Wu et al., 2024; Liu et al., 2024).
The key difference between ours and prior work is in the high-level principle: To reliably improve the translated code, we leverage the C program as the oracle in the construction of a discriminator, which automatically identifies correctness gaps and provide iterative feedback to the translator.
@@ 3.1 Code Translation Task
Given a source C program $c$, the goal of the translation is to find a functionally correct and safe Rust code $rs$ that has equivalent external behaviors as the source C program on all valid inputs. External behaviors include standard output, error messages, file modifications, and so on. More specifically, we want the translated code to satisfy the boolean requirement $\texttt{IsEq}(c,rs)$, where IsEq is an oracle that takes a C program and a Rust program as input, and determines whether they have equivalent behaviors on the valid universe $U$ of inputs that do not violate memory safety. This objective naturally captures the desired requirement that the translated Rust program has the same behavior as the C program on safe inputs. On the other hand, memory-safe handling of malicious inputs is guaranteed by the safe Rust compiler itself.
A naive approach to obtaining a Rust translation is to prompt the LLM agent with the C code $c$ and a small finite set $T\subset U$ of inputs sampled from the safe input universe $U$ with the goal of computing a Rust program that matches the behaviour of $c$ on all inputs from the sampled set $T$. Unfortunately, such a naive method does not produce a satisfactory translation. The space of valid inputs is typically infinite for most programs. Indeed, we emprically observed that with only a limited number of input instances tested one-shot by the agents, the produced translation does not generalize to unseen test cases beyond the limited set $T$.
We remark that a perfect oracle IsEq does not exist in practice. There is no automated tool that practically works in determining whether the given C code and Rust are equivalent in the input universe, or provide counterexamples when they mismatch. On the contrary, what we have is ‘IsEq*’, which takes as input both the C and the Rust code, together with a concrete input $t$, tells whether they mismatch on this input $t$. This intuitively means we run the C code and Rust code on the same input, and check if they result in the same effect. The problem now becomes how to satisfy IsEq with only IsEq* available.
@@ 3.2 Adversarial Translation
4:while NotPassesSuite($c,rs,T$) do $\triangleright$ initial translation on seed tests
@@
9:for $k=1$ to MaxIter do
10:  $numRetry\leftarrow 0$
11:  repeat
12:   $Batch\leftarrow$ Discriminator($c,rs,n=\text{TestBatchSize}$)
13:   $IsValid\leftarrow\big(\forall t\in Batch:\;\textsc{ValidTestFormat}(t)\land\textsc{SanityOk}(t,c)\big)$
14:   $IsDisc\leftarrow\big(\exists t\in Batch:\;\neg\texttt{IsEq}^{*}(c,rs,t)\big)$
15:   $numRetry\leftarrow numRetry+1$
16:   $IsReachingLimit\leftarrow(numRetry\geq\text{RetryLimit})$
17:  until $(|Batch|=\text{TestBatchSize}~\land~IsValid~\land~IsDisc)~\lor~IsReachingLimit$
18:  $T\leftarrow T\cup Batch$
19:  $numRetry\leftarrow 0$
20:  while Not PassesSuite($c,rs,T$) $\land~numRetry\leq\text{RetryLimit}$ do
21:   $numRetry\leftarrow numRetry+1$
22:   $rs\leftarrow$ Translator($c,rs,T$)
23:  end while
24:  assert IsSafe(rsrs)
25:end for
26:return rsrs as $rs^{*}$, $T$ as $T^{*}$
ACToR takes as input the C code $c$, the universe $U$, the seed test set $T_{0}$, together with hyperparameters such as the number of iterations the translator and discriminator interact (MaxIter) before ACToR outputs the final translated program. ACToR then conducts iterative translation on the source C code and outputs both the Rust translation rsrs* and the final test set TT*. During iterative improvement, ACToR maintains an append-only test set $T$. In the beginning, the translator agent generates the initial safe translation and ensures it passes all the seed tests (lines 1-3) before the adversarial iterative loop (involving the discriminator) begins. Next, in each iteration that follows, the discriminator agent reads the C code $c$ and the current Rust code $rs$, and generates test cases on which C and Rust outputs disagree (lines 7-13). The discriminator is structured so that (1) all the test cases abide by the correct format and also pass basic sanity checks when running them on the source C program (line 9); and that (2) some test cases witness a mismatch between outputs of the C and the Rust code that the translator has created so far (line 10). We ensure this by providing a prompt to the discriminator to continue retrying until the generated test cases meet the above two requirements. Once successful, the newly generated test cases ($Batch$) are added to the full test set. The translator agent then takes the updated test cases as guidance to improve the translation. It is set up with prompts that repeatedly ask it to pass all the test cases in the current test set and only use fully safe Rust (lines 16-21). If the translator runs out of retries but still cannot fully pass the test set $T$, we let it skip and move to the next iteration. After a fixed maximum number of iterations, ACToR stops the loop and outputs the final translation and the test set generated.
@@ 4 Evaluation
The goal of a good defense is to retain the functionality of the given C code, and make the defended program fully memory-safe. We evaluate our ACToR-generated Rust code on these rubrics.
Specifically, we evaluate ACToR on the following criteria:
@@
    Ablation. How much does the adversarial agentic design contribute to correctness? How do different experiment configurations influence the final results?
@@ 4.1 Experiment Settings
Our main macro-benchmark is a set of 57222The full set of BSDCoreUtils contains 69 programs, but we filter out 12 of them from our main results due to three reasons: (1) some utilities are trivial (e.g., true, yes) (2) certain utilities require special environments (e.g., stty) or privileges (e.g., sudo) not available in our experimental sandbox; and (3) certain utilities perform destructive operations (e.g., rmdir) that may break the testing harness. real-world system utility programs (e.g., ls, cp, xargs) obtained from a software project named BSDCoreUtils. It has an average of 478 lines of code each. We use this benchmark to evaluate scalability of ACToR, and compare correctness to our baselines. These programs do not come with tests or test runners. We preprocessed these programs to add test runners, which can automatically run a set of given tests for each program, compare C and Rust outputs, as well as produce coverage reports. Note that while we provide the test runner, it is infeasible to manually create comprehensive test cases for all programs to evaluate which translation strategy is more correct. Thus, unlike micro-benchmarks, we perform a relative evaluation between ACToR and baselines, as described below in the Evaluation Metrics section.
To minimize measurement noise and make our experiments automatable, our evaluation focuses on single-threaded C programs that have deterministic behaviors across different executions, which is currently an assumption of our testing framework that both the translator and the discriminator depends on. Table 1(b) summarizes the statistics of two benchmarks, where we categorize the 57 programs into five categories: process execution control, file/directory management, shell scripting commands, system information access, and textural content processing.
Agent Frameworks and Models. We use Claude Code (Anthropic, 2025a) with Claude Sonnet 4.5 (Anthropic, 2025b) as the main agent framework and LLM for our experiments. To test if our approach generalizes to other agent frameworks and LLMs, we also performed experiments using Mini-SWE-Agent (SWE-agent, 2025), a popular open-source agent framework, with GPT 5 mini (OpenAI, 2025), Claude 4 Sonnet, and Claude 4.5 Sonnet as the LLMs. We use ‘ClaudeCode+Sonnet4.5’, ‘ClaudeCode+Sonnet4‘, ‘SWE+Sonnet4.5’, ‘SWE+Sonnet4’, and ‘SWE+GPT5mini’ to represent them in later sections, respectively.
@@
Evaluation Metrics. Our primary evaluation metric to evaluate the correctness of an individual translation is pass rate, which is the percentage of tests that pass for a given set of tests. On our micro-benchmark, we report the pass rate on manually written validation tests when we evaluate different configurations of ACToR only (e.g., different LLM–agent settings or experiment hyperparameters). When we compare ACToR against other translation methods, we compute pass rates using a set of tests produced by a competing method, and refer to this metric as relative pass rate. This reduces bias from a method overfitting to its own tests (in practice, a method usually passes almost all of its own tests) and provides a stronger signal for cross-method comparison. We use relative pass rate for (i) ablation studies that compare ACToR against alternative designs on the micro-benchmark, and (ii) the final experiment on the macro-benchmark, which lacks manually written tests and therefore relies solely on test suites generated by the methods. The test sets used to compute relative pass rate achieve an average of 92.6% line coverage on the micro benchmark and 90.1% line coverage on the macro benchmark, by unioning test cases across translation methods. This indicates they are a useful signal for overall correctness as well.
Qualitative Requirements. ACToR aims to translate C code into fully safe Rust. Across both benchmarks, we enforced the absence of any unsafe blocks using a sanity check at every iteration. After all the experiments finished, we manually reviewed the final translations output by ACToR, confirming they are indeed safe and contain no malicious constructs intended to bypass the checking.
Experimental Configuration. We fixed a configuration to compare different agent settings and translation methods on the micro-benchmark. By default, ACToR and all baselines start with 15 manually crafted, diverse seed tests based on the C code for each program  [^3] The seed tests carry intended test format that both translator and discriminator of ACToR use. It also avoids a cold-start that produces translations in overly poor quality at the first few iterations for ACToR as well as all other systems we compare to.. For the coverage baseline, ACToR-NoFuzz, and ACToR, we run 10 iterations of test generation and translation. On each iteration, we generate 3 new tests, resulting in a total of 45 tests per program in the end. We fix these parameters because exploring different configurations for all agent–model settings and translation methods would be prohibitively costly. We conduct an ablation experiment to study the influence of different experiment configurations on the final correctness and cost of ACToR. We apply the most cost-effective configuration to translate the macro-benchmark. The prompt template for the translator agent is shared among all the translation methods and put in Appendix A. The prompts for the test generator agent in the coverage baseline and the discriminator of ACToR are in Appendix A.